# xDeepServe: Model-as-a-Service on Huawei CloudMatrix384

### v5

@@ Abstract
The rise of scaled-out LLMs and scaled-up SuperPods signals a new era in large-scale AI infrastructure. LLMs continue to scale out via MoE, as seen in recent models like DeepSeek, Kimi, and Qwen. In parallel, AI hardware is scaling up, with Huawei’s CloudMatrix384 SuperPod offering hundreds of GB/s high-speed interconnects. Running large MoE models on SuperPod-scale hardware brings new challenges. It requires new execution models, scalable scheduling, efficient expert load balancing, and elimination of single points of failure.
This paper presents xDeepServe, Huawei Cloud’s LLM serving system designed for SuperPod-scale infrastructure. At its core is Transformerless, a disaggregated architecture that decomposes transformer models into modular units—attention, feedforward, and MoE—executed independently on NPUs connected via high-speed fabric. We implement this design in two forms: disaggregated prefill-decode and disaggregated MoE-attention. This fully disaggregated setup enables independent scaling of compute and memory without sacrificing performance. To support this architecture, we propose XCCL, a communication library that leverages CloudMatrix384’s global shared memory to implement efficient point-to-point and all-to-all primitives. We also extend our serving engine FlowServe with system-level techniques, enabling scalable inference across hundreds of NPUs.
Figure 1: xDeepServe over CloudMatrix384 SuperPod. We deploy both disaggregated Prefill–Decode and disaggregated MoE–Attention across heterogeneous clusters. Red lines indicate requests sent to prefill, blue lines represent requests forwarded from prefill to decode.
@@ 1 Introduction
Large language models (LLMs) keep growing in size and complexity. To improve quality and efficiency, many recent models, such as DeepSeek liu2024deepseek-v3 , Kimi K2 kimi-k2 , and Qwen yang2025qwen3technicalreport , adopt the Mixture-of-Experts (MoE) architecture. MoE scales out model capacity by activating only a small subset of experts per token, reducing compute cost while keeping high model quality. At the same time, AI hardware is scaling up. Huawei’s CloudMatrix384 connects 384 Ascend chips with a hundreds of GB/s high-speed interconnect, offering a tightly coupled memory and compute environment across the entire pod.
@@
This paper presents xDeepServe, Huawei Cloud’s in-house LLM serving system for SuperPod-scale deployment. We advocate resource disaggregation legoos-osdi18 as a foundational design principle for large-scale serving. Disaggregation decouples system components into independently scalable units, improves fault isolation, and enables flexible system evolution. To realize these benefits, we introduce Transformerless, an architecture that breaks transformer-based LLMs into modular components—attention, feedforward, and MoE—and runs each module independently on dedicated NPUs. This design decouples compute and memory, reduces interference across stages, and supports separate scaling.
We integrate two forms of disaggregation.
@@
First, to reduce interference between compute-bound prefill and memory-bound decode phases, we use disaggregated prefill-decode architecture (§5.1). While prior systems like Splitwise patel2023splitwise , TetriServe tetriserve-arxiv24 , and DistServe zhong2024distserve explored this direction, our design targets the unique demands of large MoE models with expert parallelism on SuperPod-scale hardware. This disaggregated pipeline is further enhanced by heterogeneous deployment: we run prefill on both Ascend 910 (previous generation) and CloudMatrix384 NPUs, while decode runs exclusively on CloudMatrix’s nodes to leverage high-bandwidth interconnects. This design balances cost and performance, while achieving high throughput and low latency at scale.
@@
Second, we further disaggregate MoE and attention, running them on separate NPUs to reduce resource contention and improve utilization (§5.2). While prior work explored this idea on smaller models, xDeepServe is the first to apply it at SuperPod scale, deploying DeepSeek-V3/R1 across 768 Ascend NPU dies, with 288 handling MoE and 480 handling attention. This separation introduces a compute imbalance: MoE is stateless and scales with batch size, while attention maintains KV cache and scales with sequence length. To address this, we introduce three techniques: (1) a trampoline-based A2E/E2A routing protocol to handle asymmetric NPU allocation, (2) a new DP domain abstraction to enable inter-group scheduling without shrinking batch sizes, and (3) persistent kernel scheduling across concurrent compute and communication streams to eliminate CPU overhead. These designs allow us to fully utilize both MoE and attention NPUs while maintaining sub-200 µs dispatch latency across the entire SuperPod.
@@
To support fully disaggregated execution, we propose XCCL, a novel communication library that provides memory-semantic point-to-point and all-to-all primitives over CloudMatrix384’s shared memory fabric. XCCL supports microsecond-level latency and high concurrency across more than 300,000 NPU pairs. We use it to build KV cache transfer, MoE dispatch/combine, and MoE-attention (§3).
@@
xDeepServe further builds a scalable serving engine at SuperPod-Scale called FlowServe. It introduces the Data Parallel (DP) group abstraction, where each group manages its own execution pipeline—including tokenization, SPMD execution, caching, and networking—without relying on central coordination. FlowServe scales to hundreds of NPUs by eliminating single points of scalability bottlenecks and applying system-level techniques to reduce latency under MoE barriers such as expert dispatch and combine. These include load-aware request routing among DPs, proactive garbage collection to reduce jitter, and MoE expert load balancing. To improve efficiency and hardware utilization, FlowServe supports Multi-Token Prediction (MTP) and INT8 quantization, enabling high-throughput inference under tight SLA constraints.
@@
xDeepServe is deployed in production to serve large-scale DeepSeek liu2024deepseek-v3 , Kimi kimi-k2 , and Qwen yang2025qwen3technicalreport models. It achieves 2400 tokens per second per Ascend NPU chip while meeting a 50s time-per-output-token (TPOT) SLA. In the rest of the paper, we first describe the architecture of CloudMatrix384 in §2.2, followed by the low-level communication library in §3. We then present the serving engine in §4 and the Transformerless serving architecture in §5. Finally, we discuss our reliability mechanisms in §6.
@@ 2 Background
xDeepServe is Huawei Cloud’s fully-hosted, serverless LLM serving platform, designed to support large-scale, multi-tenant workloads. It evolves from the DeepServe system introduced in our previous work hu2025deepflow , with new capabilities targeting LLM serving at the scale of CloudMatrix384. xDeepServe has been in production for over a year, operating on a large Ascend NPU cluster. It provides industry-standard APIs for fine-tuning, agent hosting, and model serving. Motivated by the growing demand for generative AI services (e.g., ChatGPT and DeepSeek), xDeepServe addresses key production challenges, including heterogeneous workload durations, distributed stateful execution, and dynamic resource demands.
To tackle these challenges, xDeepServe introduces four core components hu2025deepflow . First, it adopts a serverless abstraction based on a request–job–task model to manage diverse AI workloads, spanning post-training and inference tasks. Second, it integrates a custom serving engine, FlowServe, which follows a microkernel-inspired architecture, supports NPU-centric execution, and uses SPMD-style parallelism for high-performance LLM serving. Third, it includes scheduling policies designed to support both prefill-decode (PD) disaggregated and PD-colocated deployment modes. Fourth, it incorporates system-level optimizations, such as pre-warmed pods, DRAM preloading, and NPU fork, to enable rapid elasticity, scaling to 64 instances within seconds.
While our previous paper hu2025deepflow introduced the core architecture of xDeepServe, this report focuses on the new challenges and techniques arising from deployment on the CloudMatrix384 SuperPod. We present our latest system designs and methods for leveraging its scaled-up capabilities to enable next-generation LLM serving.
@@ 2.2 CloudMatrix384
Figure 2: An Overview of CloudMatrix384 SuperPod. A CloudMatrix384 SuperPod has 48 servers and 384 Ascend NPU chips in total. A single CloudMatrix384 Ascend NPU chip has two dies interconnected via high-bandwidth NoC. The chip uses the decoupled DaVinci architecture.
Figure 2 shows an overview of CloudMatrix384. A CloudMatrix384 SuperPod is a single scale-up domain composed of 48 servers, delivering hundreds of PFLOPs of FP16 compute, several terabytes of on-chip memory, and terabytes per second of memory bandwidth. Each server in the SuperPod is equipped with multiple CPUs, multiple NICs, and 8 Ascend NPU chips. Crucially, the SuperPod integrates three types of network fabrics. The first is a traditional VPC network that connects servers to external systems or cloud services. The second is a scaled-out RoCE network that interconnects all NPU chips and can be extended across multiple SuperPods and previous generation Ascend NPU servers. The third is a scaled-up UB network that makes CloudMatrix384 a SuperPod. It connects all CPUs and NPUs within the SuperPod in an all-to-all fashion, offering roughly several times the bandwidth of RoCE. See  ascend-npu-hotchips19 ; ascend-c-programming-manual ; cloudmatrix384 for more.
@@ 2.3 NPU Execution Mode
Single-op Mode. This is PyTorch’s default execution mode. Each operator is dispatched to the NPU operator queue upon Python invocation, and its result is accessible on next synchronization. This dynamic style supports arbitrary control flow, dynamic tensor shapes, and native debugging, and it allows limited CPU–NPU concurrency when dependencies permit. However, it introduces overhead because each operator launch incurs host-device communication, and the NPU may stay idle if the operator execution time is lower than the dispatch time of next operator. That said, for computation-heavy operators, this cost is acceptable. Therefore, we use this mode during prefill to handle dynamic input shapes.
Graph Mode. Graph mode traces and compiles a static computation graph, which is then dispatched to the NPU in a single kernel launch. This eliminates per-operator launch overhead and improves device utilization. Huawei’s TorchAir (based on PTA) enables graph-mode inference on Ascend NPUs. As shown in Figure 3(b), the CPU submits the entire computation graph in one operation, reducing dispatch overhead. We use this mode during decode, where inputs are small and mostly static, to achieve optimal performance.
@@ 3 Communication Library over CloudMatrix’s Distributed Shared Memory
A high-performance communication library is essential to fully harness the scaled-up UB fabric in CloudMatrix384. We introduce XCCL, a communication library purpose-built for LLM serving on SuperPod. XCCL proposes distributed memory protocols over CloudMatrix384’s global shared memory to construct efficient networking primitives. Its protocol design resembles classical far-memory systems based on one-siden RDMA verbs farm-nsdi14 ; clover-atc20 . Note that the memory transfer units (mtu) in NPU’s computing cores, along with the DMA engines, are fundamental to enabling global shared memory in CloudMatrix384. Both can read from or write to the on-chip memory of any other NPU in the SuperPod via the UB fabric. This capability underpins our customized communication library for serving. In general, memory transfer units are used for low-latency communication, while DMA engines are preferred for high-throughput transfers. This is because memory transfer units provide a memory-semantic API and move data between the computing core’s data buffer and on-chip memory, with transfers limited to the buffer size of hundreds of KBs. In contrast, DMA engines use DMA semantics and support bulk data movement up to several GBs.
Below, we describe the point-to-point APIs used in disaggregated prefill and decode (§3.1), the dispatch and combine APIs for expert parallelism (EP) (§3.2), and the A2E/E2A APIs for disaggregated MoE and attention (§3.3).
@@ 3.1 Point-to-Point Primitives
Figure 4: Distributed Send/Recv Workflow. We only show memory-semantic-based transfer using memory transfer units while memory copy between on-chip memory can also be done by using the DMA engine. We also have a zero-copy version in which the send and recv kernel directly manipulate the app data area.
@@
This section describes the design and implementation of point-to-point APIs such as send and receive on the CloudMatrix384 SuperPod. These APIs are essential for serving, enabling efficient data transfer in scenarios such as disaggregated prefill and decode, sequence parallelism, and fast scaling via npu-fork hu2025deepflow . For simplicity, we use the transfer of KV cache hu2025epic from prefill NPU to decode NPU to illustrate the send/receive design. XCCL’s send and receive operations can occur between any pair of NPUs in CloudMatrix384, hence our design can scale to roughly 300K potential pairs.
@@
Data structure. Each NPU’s on-chip memory is partitioned into three areas: the app data area, the metadata area, and the managed data area. The app data area is reserved for application-specific data such as KV cache, used directly by our serving engine. The metadata area stores control information required for sanity checks and to implement the distributed memory protocol. It consists of metadata fields, one for each pair of computing cores to enable parallelism among computing cores. Given that CloudMatrix384 includes 384 NPU chips, each with 2 dies and up to 48 computing cores per die, the total number of fields is $384\times 2\times 48\times 2\approx 74\text{K}$. Each 32-byte field includes a user-defined eventID for sanity checking, a kernel-generated chunkID for tracking chunked transfers, and a tailPtr pointing to the ring buffer in the data area. The total metadata size is set to 4 MB. The managed data area is used for data exchange between NPUs. A dedicated ring buffer is maintained for each NPU pair, with a fixed number of slots of fixed size per buffer. Both the metadata and managed data areas are managed by XCCL. The current implementation is not zero-copy; data is copied between the app data area and the managed data area when XCCL APIs are invoked.
@@
Distributed Memory Protocol. We design an distributed protocol to transfer data from sender to receiver over CloudMatrix384’s global shared memory, illustrated in Figure 4. Step 1: The sender’s serving engine invokes XCCL’s send, passing the source buffer in the app data area (e.g., KV cache), an eventID (e.g., number of sends), the receiver NPU’s ID, and the number of computing cores to use. XCCL launches a kernel on the sender NPU. The send kernel uses mtu-in to copy data from the app data area to each computing core’s data buffer in parallel. Step 2: The send kernel then reads the destination’s on-chip memory address from its metadata area and uses mtu-out to transfer data from the data buffer to the receiver’s managed data area. Each data buffer operates in a ping-pong fashion, allowing mtu-in and mtu-out to run concurrently. Alternatively, the send kernel can also directly copy data from the app data area to the receiver’s managed data area directly using the DMA engine, which bypass the data buffer but at the cost of higher starting latency. Step 3: The send kernel updates the receiver’s tailPtr in the metadata area via mtu-out, indicating the amount of data transferred. Step 4: The send kernel then busy-polls its local metadata area for an acknowledgment from the receiver. Step 5: Meanwhile, the receiver’s serving engine invokes XCCL’s receive, passing the destination buffer in the app data area, the eventID, the sender NPU’s ID, and the number of computing cores. XCCL launches a kernel on the receiver NPU, which polls the metadata area for new data. Step 6: Upon detecting data, the receive kernel copies it from the managed data area to the app data area using mtu-in and mtu-out, again in a ping-pong fashion. Step 7: After the transfer completes, the receive kernel updates the sender’s metadata area to acknowledge completion, then returns control to the CPU and the receive caller. Step 8: The send kernel detects the acknowledgment, returns control to the CPU, and completes the send call. In addition to this synchronous protocol, XCCL supports an asynchronous mode in which send and receive kernels avoid busy polling.
@@
Figure 5: Evaluation of Send/Recv. We vary the data size and the number of computing cores used for a single send/recv pair.
Figure 6: Evaluation of Dispatch/Combine. We vary the batch size per die with a fixed EP128. We use DeepSeek-R1.

### v6

# Huawei Cloud Model-as-a-Service on the CloudMatrix384 SuperPod

@@ Abstract
Abstract. Scaled-out MoE LLMs and scaled-up SuperPods create new systems challenges for production Model-as-a-Service (MaaS), requiring disaggregation, low-latency communication, and decentralized serving. This report presents xDeepServe, the production serving system behind Huawei Cloud’s MaaS offering on CloudMatrix384, a 48-server SuperPod with 384 Ascend 910C chips connected by a high-bandwidth UB fabric and global shared memory. It serves models including DeepSeek, Kimi, GLM, Qwen, and MiniMax, among others.
xDeepServe is built around Transformerless, a disaggregated execution architecture that decomposes transformer inference into modular units—attention, feedforward, and MoE—and supports disaggregated Prefill-Decode and MoE-Attention deployments. To enable disaggregation, we develop XCCL, a memory-semantic communication layer providing microsecond-level point-to-point and scalable all-to-all primitives, and we extend FlowServe with decentralized DP groups and techniques to mitigate stragglers and synchronization variance. In a peak decoding configuration, xDeepServe reaches 2400 tokens/s per Ascend 910C chip at $\sim$50 ms time-per-output-token (TPOT).
Figure 1: xDeepServe Architecture over CloudMatrix384 SuperPod.
Contents
@@ 1 Introduction
Large language models (LLMs) keep growing in size and complexity. To improve quality and efficiency, many recent models, such as DeepSeek [liu2024deepseek-v3], Kimi K2 [kimi-k2], and Qwen [yang2025qwen3technicalreport], adopt the Mixture-of-Experts (MoE) architecture. MoE scales out model capacity by activating only a small subset of experts per token, reducing compute cost while keeping high model quality. At the same time, AI hardware is scaling up. Huawei’s CloudMatrix384 connects 384 Ascend 910C chips with hundreds of GB/s of high-speed interconnect bandwidth, offering a tightly coupled memory and compute environment across the entire pod.
@@
This paper presents xDeepServe, Huawei Cloud’s in-house LLM serving system for SuperPod-scale deployment.
@@
We advocate resource disaggregation [legoos-osdi18] as a foundational design principle for large-scale serving. Disaggregation decouples system components into independently scalable units, improves fault isolation, and enables flexible system evolution. To realize these benefits, we introduce Transformerless, an architecture that breaks transformer-based LLMs into modular components—attention, feedforward, and MoE—and runs each module independently on dedicated NPUs. This design decouples compute and memory, reduces interference across stages, and supports separate scaling.
We implement two forms of disaggregation.
@@
First, to reduce interference between compute-bound prefill and memory-bound decode phases, we use a disaggregated Prefill-Decode architecture (§5.1). While prior systems like Splitwise [patel2023splitwise], TetriServe [tetriserve-arxiv24], and DistServe [zhong2024distserve] explored this direction, our design targets the unique demands of large MoE models with expert parallelism on SuperPod-scale hardware. This disaggregated pipeline is further enhanced by heterogeneous deployment: we run prefill on both Ascend 910B and 910C NPUs, while decode runs exclusively on CloudMatrix’s 910C nodes to leverage high-bandwidth interconnects. This design balances cost and performance, while achieving high throughput and low latency at scale.
@@
Second, we further disaggregate MoE-Attention by running MoE experts and attention on separate NPUs to reduce resource contention and improve utilization (§5.2). While prior work explored this idea on smaller models, we demonstrate this design at SuperPod scale by deploying DeepSeek-V3/R1 across 768 Ascend 910C dies, with 288 handling MoE and 480 handling attention. This separation introduces a compute imbalance: MoE is stateless and scales with batch size, while attention maintains KV cache and scales with sequence length. To address this, we introduce three techniques: (1) a trampoline-based A2E/E2A routing protocol to handle asymmetric NPU allocation, (2) a new DP domain abstraction to enable inter-group scheduling without shrinking batch sizes, and (3) persistent kernel scheduling across concurrent compute and communication streams to eliminate CPU overhead. These designs allow us to fully utilize both MoE and attention NPUs while maintaining sub-200 $\mu s$dispatch latency across the entire SuperPod.
@@
To support fully disaggregated execution, we propose XCCL, a novel communication library that provides memory-semantic point-to-point and all-to-all primitives over CloudMatrix384’s shared-memory fabric. XCCL supports microsecond-level latency and high concurrency across more than 300,000 NPU pairs. We use it to build KV cache transfer, MoE dispatch/combine, and MoE-Attention (§3).
@@
xDeepServe further builds a scalable serving engine at SuperPod scale, called FlowServe. It introduces the Data Parallel (DP) group abstraction, where each group manages its own execution pipeline—including tokenization, SPMD execution, caching, and networking—without relying on central coordination. FlowServe scales to hundreds of NPUs by eliminating single points of scalability bottlenecks and applying system-level techniques to reduce latency under MoE barriers such as expert dispatch and combine. These include load-aware request routing among DPs, proactive garbage collection to reduce jitter, and MoE expert load balancing. To improve efficiency and hardware utilization, FlowServe supports Multi-Token Prediction (MTP) and INT8 quantization, enabling high-throughput inference under tight SLA constraints.
@@
xDeepServe is deployed in production to serve large-scale DeepSeek [liu2024deepseek-v3], GLM [glm5team2026glm5vibecodingagentic], Kimi [kimi-k2], and Qwen [yang2025qwen3technicalreport] models. In our peak decoding configuration, it achieves 2400 tokens/s per Ascend 910C chip at $\sim$50 ms time-per-output-token (TPOT). In production, we target a TTFT SLA under 2 s and a TPOT SLA of 35 ms in most cases.
Contributions. We make the following contributions:
  • We present xDeepServe, a production MaaS serving system for running large MoE LLMs on CloudMatrix384 SuperPods.
  • We describe Transformerless, including disaggregated Prefill-Decode and disaggregated MoE-Attention deployments at SuperPod scale.
  • We introduce XCCL, a communication layer over CloudMatrix384’s distributed shared memory that provides microsecond-level point-to-point and all-to-all primitives for serving.
  • We present the SuperPod-scale serving design of FlowServe, including decentralized DP groups, load-aware scheduling, expert load balancing, and reliability mechanisms.
In the rest of the paper, we first describe the architecture of CloudMatrix384 in §2.2, followed by the low-level communication library in §3. We then present the serving engine in §4 and the Transformerless serving architecture in §5. Finally, we discuss our reliability mechanisms in §6.
@@ 2 Background
xDeepServe is Huawei Cloud’s fully-hosted, serverless LLM serving platform, designed to support large-scale, multi-tenant workloads. It evolves from the DeepServe system introduced in our previous work [hu2025deepflow], with new capabilities targeting LLM serving at the scale of CloudMatrix384. xDeepServe has been in production for over a year, operating on a large Ascend NPU cluster. It provides industry-standard APIs for fine-tuning, agent hosting, and model serving. Motivated by the growing demand for generative AI services (e.g., ChatGPT and DeepSeek), xDeepServe addresses key production challenges, including heterogeneous workload durations, distributed stateful execution, and dynamic resource demands.
To tackle these challenges, xDeepServe introduces four core components [hu2025deepflow]. First, it adopts a serverless abstraction based on a request-job-task model to manage diverse AI workloads, spanning post-training and inference tasks. Second, it integrates a custom serving engine, FlowServe, which follows a microkernel-inspired architecture, supports NPU-centric execution, and uses SPMD-style parallelism for high-performance LLM serving. Third, it includes scheduling policies designed to support both Prefill-Decode (PD) disaggregated and PD-colocated deployment modes. Fourth, it incorporates system-level optimizations, such as pre-warmed pods, DRAM preloading, and NPU fork, to enable rapid elasticity, scaling to 64 instances within seconds.
While our previous paper [hu2025deepflow] introduced the core architecture of xDeepServe, this report focuses on the new challenges and techniques arising from deployment on the CloudMatrix384 SuperPod. We present our latest system designs and methods for leveraging its scaled-up capabilities to enable next-generation LLM serving.
@@ 2.2 CloudMatrix384
Figure 2: An Overview of CloudMatrix384 SuperPod and Ascend 910C NPU Chip. A CloudMatrix384 SuperPod has 48 servers and 384 Ascend 910C NPU chips in total. A single 910C NPU chip has two dies interconnected via high-bandwidth NoC. The chip uses the decoupled DaVinci architecture, including independent AI cube, AI vector, etc.
Figure 2 shows an overview of CloudMatrix384 and the 910C NPU chip.
A CloudMatrix384 SuperPod is a single scale-up domain composed of 48 servers, delivering hundreds of PFLOPs of FP16 compute, several terabytes of on-chip memory, and terabytes per second of memory bandwidth. Each server in the SuperPod is equipped with multiple CPUs, multiple NICs, and 8 Ascend 910C NPU chips. Crucially, the SuperPod integrates three types of network fabrics. The first is a traditional VPC network that connects servers to external systems or cloud services. The second is a scaled-out RoCE network that interconnects all NPU chips and can be extended across multiple SuperPods and 910B servers. The third is a scaled-up UB network that makes CloudMatrix384 a SuperPod. It connects all CPUs and NPUs within the SuperPod in an all-to-all fashion, offering several times higher bandwidth than RoCE.
@@
An Ascend 910C NPU chip consists of two dies interconnected via an on-chip network. Each die follows the DaVinci architecture [ascend-npu-hotchips19, ascend-c-programming-manual, cloudmatrix384], comprising an equal number of AI Cube (AIC) and AI Vector (AIV) cores, several AI CPUs, and multiple DMA engines. AIC cores execute matrix-oriented kernels, while AIV cores handle vector operations. Both core types include scalar units and support independent control flow, with data exchange via external on-chip memory. Each AIV core includes a KB-level unified buffer and two Memory Transfer Engines (MTEs), codenamed MTE2 and MTE3, which transfer data between on-chip memory and the unified buffer: MTE2 loads data from on-chip memory into the buffer, and MTE3 writes data back to on-chip memory.
The MTE2/MTE3 engines in AIV cores, along with the DMA engines, are fundamental to enabling global shared memory in CloudMatrix384. Both MTE2/MTE3 and DMA engines on an NPU can read from or write to the on-chip memory of any other NPU in the SuperPod via the UB fabric. This capability underpins our customized communication library for serving (§3). In general, MTE2/MTE3 are used for low-latency communication, while DMA engines are preferred for high-throughput transfers. This is because MTE2/MTE3 provide a memory-semantic API and move data between the AIV’s unified buffer and on-chip memory, with transfers limited to the buffer size of hundreds of KBs. In contrast, DMA engines use DMA semantics and support bulk data movement up to several GBs. For more details, refer to the Ascend C Programming Manual.
@@ 2.3 NPU Execution Mode
  • Single-op Mode. This is PyTorch’s default execution mode. Each operator is dispatched to the NPU operator queue upon Python invocation, and its result is accessible on the next synchronization. This dynamic style supports arbitrary control flow, dynamic tensor shapes, and native debugging, and it allows limited CPU-NPU concurrency when dependencies permit. However, it introduces overhead because each operator launch incurs host-device communication, and the NPU may stay idle if the operator execution time is lower than the dispatch time of the next operator. That said, for computation-heavy operators, this cost is acceptable. Therefore, we use this mode during prefill to handle dynamic input shapes.
  • Graph Mode. Graph mode traces and compiles a static computation graph, which is dispatched as a single unit to the NPU. This removes per-operator launch overhead and maximizes device utilization. Huawei’s TorchAir (built on PTA) supports graph-mode inference on Ascend NPUs. As shown in Figure 3(b), the CPU can dispatch the entire graph to the NPU in one kernel launch. We use graph mode during decode for optimal performance, as input shapes are small and mostly static.
@@ 3 Communication Library over CloudMatrix’s Distributed Shared Memory
A high-performance communication library is essential to fully harness the scaled-up UB fabric in CloudMatrix384. We introduce XCCL, a communication library purpose-built for LLM serving on SuperPod. XCCL proposes distributed memory protocols over CloudMatrix384’s global shared memory to construct efficient networking primitives. Its protocol design resembles classical far-memory systems based on one-sided RDMA verbs [farm-nsdi14, clover-atc20].
We describe the point-to-point APIs used in disaggregated Prefill-Decode (§3.1), the dispatch and combine APIs for expert parallelism (EP) (§3.2), and the A2E/E2A APIs for disaggregated MoE-Attention (§3.3).
@@ 3.1 Point-to-Point Primitives
Figure 4: Distributed Send/Receive Workflow. We show two NPUs and the distributed memory transfer protocol. We only show memory-semantic-based transfer using MTE2/MTE3 while remote memory copies can also be performed using the DMA engine. We also have a zero-copy version in which the send and receive kernels directly manipulate the app data area.
@@
This section describes the design and implementation of point-to-point APIs such as send and receive on the CloudMatrix384 SuperPod. These APIs are essential for serving, enabling efficient data transfer in scenarios such as disaggregated prefill and decode, sequence parallelism, and fast scaling via npu-fork [hu2025deepflow]. For simplicity, we use the transfer of KV cache [hu2025epic] from prefill NPU to decode NPU to illustrate the send/receive design. XCCL’s send and receive operations can occur between any pair of NPUs in CloudMatrix384, hence our design can scale to roughly 300K potential pairs.
We now describe the design details.
@@
Data structure. Each NPU’s on-chip memory is partitioned into three areas: the app data area, the metadata area, and the managed data area. The app data area is reserved for application-specific data such as KV cache, used directly by our serving engine. The metadata area stores control information required for sanity checks and to implement the distributed memory protocol. It consists of metadata fields, one for each pair of AIV cores to enable parallelism among AIV cores. Given that CloudMatrix384 includes 384 NPU chips, each with 2 dies and up to 48 AIV cores per die, the total number of fields is $384\times 2\times 48\times 2\approx 74\text{K}$. Each 32-byte field includes a user-defined eventID for sanity checking, a kernel-generated chunkID for tracking chunked transfers, and a tailPtr pointing to the ring buffer in the data area. The total metadata size is set to 4 MB. The managed data area is used for data exchange between NPUs. A dedicated ring buffer is maintained for each NPU pair, with a fixed number of slots of fixed size per buffer. Both the metadata and managed data areas are managed by XCCL. The current implementation is not zero-copy; data is copied between the app data area and the managed data area when XCCL APIs are invoked.
@@
Distributed Memory Protocol. We design a distributed protocol to transfer data from sender to receiver over CloudMatrix384’s global shared memory, illustrated in Figure 4. Step 1: The sender’s serving engine invokes XCCL’s send, passing the source buffer in the app data area (e.g., KV cache), an eventID (e.g., number of sends), the receiver NPU’s ID, and the number of AIV cores to use. XCCL launches a kernel on the sender NPU. The send kernel uses MTE2 to copy data from the app data area to each AIV’s unified buffer in parallel. Step 2: The send kernel then reads the destination’s on-chip memory address from its metadata area and uses MTE3 to transfer data from the unified buffer to the receiver’s managed data area. Each unified buffer operates in a ping-pong fashion, allowing MTE2 and MTE3 to run concurrently. Alternatively, the send kernel can copy data from the app data area to the receiver’s managed data area using the DMA engine, which bypasses the unified buffer but incurs higher startup latency. Step 3: The send kernel updates the receiver’s tailPtr in the metadata area via MTE3, indicating the amount of data transferred. Step 4: The send kernel then busy-polls its local metadata area for an acknowledgment from the receiver. Step 5: Meanwhile, the receiver’s serving engine invokes XCCL’s receive, passing the destination buffer in the app data area, the eventID, the sender NPU’s ID, and the number of AIV cores. XCCL launches a kernel on the receiver NPU, which polls the metadata area for new data. Step 6: Upon detecting data, the receive kernel copies it from the managed data area to the app data area using MTE2 and MTE3, again in a ping-pong fashion. Step 7: After the transfer completes, the receive kernel updates the sender’s metadata area to acknowledge completion, then returns control to the CPU and the receive caller. Step 8: The send kernel detects the acknowledgment, returns control to the CPU, and completes the send call. In addition to this synchronous protocol, XCCL supports an asynchronous mode in which send and receive kernels avoid busy polling.
@@
Figure 5: Evaluation of Send/Receive. We vary the data size and the number of AIV cores used for a single send/receive pair.
Figure 6: Evaluation of Dispatch/Combine. We vary the batch size per die with a fixed EP128. Dispatch includes a quantization step.